What is the perimeter of a rectangle whose width is (x-3)and length is (3y+5)? What is the the area of that rectangle?
Answer by JoelSchwartz(130)

w=x-3
L=3y+5
p=2w+2L
p=2(x-3)+2(3y+5)
p=2x-6+6y+10
p=2x+6x+4
A=wL
A=(x-3)(3y+5)
area of this rectangle
A=3yx-9y+5x-15